# EgoExoLearn: A Dataset for Bridging Asynchronous Ego- and Exo-centric View of Procedural Activities in Real World

### v1

@@ Abstract
Being able to map the activities of others into one’s own point of view is a fundamental human skill even from a very early age. Taking a step toward understanding this human ability, we introduce EgoExoLearn, a large-scale dataset that emulates the human demonstration following process, in which individuals record egocentric videos as they execute tasks guided by exocentric-view demonstration videos. Focusing on the potential applications in daily assistance and professional support, EgoExoLearn contains egocentric and demonstration video data spanning 120 hours captured in daily life scenarios and specialized laboratories. Along with the videos we record high-quality gaze data and provide detailed multimodal annotations, formulating a playground for modeling the human ability to bridge asynchronous procedural actions from different viewpoints. To this end, we present benchmarks such as cross-view association, cross-view action planning, and cross-view referenced skill assessment, along with detailed analysis. We expect EgoExoLearn can serve as an important resource for bridging the actions across views, thus paving the way for creating AI agents capable of seamlessly learning by observing humans in the real world. The dataset and benchmark codes are available at https://github.com/OpenGVLab/EgoExoLearn/.
footnotetext: $*{}^{*}$orresponding authors. ${}^{\dagger}$roject lead. ${}^{\ddagger}$qual key contributions.
@@ 3.2 Annotation
Coarse-level language annotation. In this step, we ask annotators to annotate the coarse actions in the videos. Like the previous works [106, 126], the coarse actions are defined as a middle-level step for accomplishing a task and can be divided into multiple fine actions. For instance, “Prepare the pork” in the task of twice-cooked pork, and “Suction filtration” in solid-phase peptide synthesis. Three types of annotation are given in this step: 1) the temporal interval, consisting of start and end timestamps; 2) the action label; and 3) a language description of the video within the annotated interval. We specifically request annotators to focus on elucidating “what is done,” “how it is done,” and “the purpose of this step” in their language descriptions. We define a total of 39 categories of coarse-level actions in this stage and acquire 41.2 coarse-level annotations per video with an average length of 21.5 seconds.
Fine-level language annotation. Based on the coarse-level annotations, in this step we request the annotators to provide annotations for the fine-level actions. The fine-level actions are the atomic actions like “take knife” or “pull syringe plunger”. Unlike the first step, annotators are instructed to furnish language descriptions that specifically emphasize “which hand is used”, “what object is used” and “why it is used”. For the first two steps, we employ a two-round manual annotation checking to ensure the annotation quality.
@@
Table 1: Comparison to related datasets on settings (left) and annotations (right). “Unique hours” refers to the cumulative duration of distinct video recordings, counting only one camera’s footage for simultaneous recordings of the same activity. normal-$\diamond$: Partially included.
@@ 3.3 Statistics & Comparisons
To the best of our knowledge, there is no dataset that follows the same setting as ours for a direct comparison. Therefore, we enumerate various aspects of our dataset and conduct a comparative analysis with relevant datasets in Tables 1 and 2. EgoExoLearn distinctively enriches the domain with its “visual demonstration following” setting. Beyond this unique setting, it stands as the first egocentric dataset that includes temporal bounded language captions, annotated cross-view associations, and multi-label video segments.
@@ 4.2.2 Cross-view action anticipation & planning
We comprehensively consider four evaluation settings. The “Ego-only” and “Exo-only” settings do not involve cross-view understanding, thus we use zero-shot evaluation serving as the references. The Ego2Exo and Exo2Ego settings are the cross-view settings. For UDA, “Ego2Exo” is defined as utilizing the egocentric view as the source domain and the exocentric view as the target domain. In the context of KD, “Ego2Exo” indicates we initially train a teacher model on egocentric data, followed by the training and distillation of the student model on exocentric data. For CT, we merge both egocentric and exocentric datasets through direct concatenation. We report results on the test set and put the validation set results in the supplementary.

### v2

@@ Abstract
Being able to map the activities of others into one’s own point of view is a fundamental human skill even from a very early age. Taking a step toward understanding this human ability, we introduce EgoExoLearn, a large-scale dataset that emulates the human demonstration following process, in which individuals record egocentric videos as they execute tasks guided by exocentric-view demonstration videos. Focusing on the potential applications in daily assistance and professional support, EgoExoLearn contains egocentric and demonstration video data spanning 120 hours captured in daily life scenarios and specialized laboratories. Along with the videos we record high-quality gaze data and provide detailed multimodal annotations, formulating a playground for modeling the human ability to bridge asynchronous procedural actions from different viewpoints. To this end, we present benchmarks such as cross-view association, cross-view action planning, and cross-view referenced skill assessment, along with detailed analysis. We expect EgoExoLearn can serve as an important resource for bridging the actions across views, thus paving the way for creating AI agents capable of seamlessly learning by observing humans in the real world. The dataset and benchmark codes are available at https://github.com/OpenGVLab/EgoExoLearn.
footnotetext: Corresponding authors. Project lead. Equal key contributions.
@@ 3.2 Annotation
Coarse-level language annotation. In this step, we ask annotators to annotate the coarse actions in the videos. Like the previous works [116, 137], the coarse actions are defined as a middle-level step for accomplishing a task and can be divided into multiple fine actions. For instance, “Prepare the pork" in the task of twice-cooked pork, and “Suction filtration" in solid-phase peptide synthesis. Three types of annotation are given in this step: 1) the temporal interval, consisting of start and end timestamps; 2) the action label; and 3) a language description of the video within the annotated interval. We specifically request annotators to focus on elucidating “what is done," “how it is done," and “the purpose of this step" in their language descriptions. We define a total of 39 categories of coarse-level actions in this stage and acquire 41.2 coarse-level annotations per video with an average length of 21.5 seconds.
Fine-level language annotation. Based on the coarse-level annotations, in this step we request the annotators to provide annotations for the fine-level actions. The fine-level actions are the atomic actions like “take knife" or “pull syringe plunger". Unlike the first step, annotators are instructed to furnish language descriptions that specifically emphasize “which hand is used", “what object is used" and “why it is used". For the first two steps, we employ a two-round manual annotation checking to ensure the annotation quality.
@@
Table 1: Comparison to related datasets on settings (left) and annotations (right). “Unique hours" refers to the cumulative duration of distinct video recordings, counting only one camera’s footage for simultaneous recordings of the same activity. $\diamond$: Partially included.
@@ 3.3 Statistics & Comparisons
To the best of our knowledge, there is no dataset that follows the same setting as ours for a direct comparison. Therefore, we enumerate various aspects of our dataset and conduct a comparative analysis with relevant datasets in Tables 1 and 2. EgoExoLearn distinctively enriches the domain with its “visual demonstration following" setting. Beyond this unique setting, it stands as the first egocentric dataset that includes temporal bounded language captions, annotated cross-view associations, and multi-label video segments.
@@ 4.2.2 Cross-view action anticipation & planning
We comprehensively consider four evaluation settings. The “Ego-only" and “Exo-only" settings do not involve cross-view understanding, thus we use zero-shot evaluation serving as the references. The Ego2Exo and Exo2Ego settings are the cross-view settings. For UDA, “Ego2Exo” is defined as utilizing the egocentric view as the source domain and the exocentric view as the target domain. In the context of KD, “Ego2Exo” indicates we initially train a teacher model on egocentric data, followed by the training and distillation of the student model on exocentric data. For CT, we merge both egocentric and exocentric datasets through direct concatenation. We report results on the test set and put the validation set results in the supplementary.